# Membership Inference on Text-to-image Diffusion Models via Conditional Likelihood Discrepancy

### v2

@@ Abstract
Text-to-image diffusion models have achieved tremendous success in the field of controllable image generation, while also coming along with issues of privacy leakage and data copyrights. Membership inference arises in these contexts as a potential auditing method for detecting unauthorized data usage. While some efforts have been made on diffusion models, they are not applicable to text-to-image diffusion models due to the high computation overhead and enhanced generalization capabilities. In this paper, we first identify a conditional overfitting phenomenon in text-to-image diffusion models, indicating that these models tend to overfit the conditional distribution of images given the corresponding text rather than the marginal distribution of images only. Based on the observation, we derive an analytical indicator, namely Conditional Likelihood Discrepancy (CLiD), to perform membership inference, which reduces the stochasticity in estimating memorization of individual samples. Experimental results demonstrate that our method significantly outperforms previous methods across various data distributions and scales. Additionally, our method shows superior resistance to overfitting mitigation strategies such as early stopping and data augmentation.
@@ 1 Introduction
Text-to-image diffusion models have achieved remarkable success in the guided generation of diverse, high-quality images based on text prompts, such as Stable Diffusion [40, 36], DALLE-2 [37], Imagen [43], and DeepFloyd-IF [26]. These models are increasingly adopted by users to create photorealistic images that align with desired semantics. Moreover, they can generate images of specific concepts [27] or styles [53] when fine-tuned on relevant datasets. However, the impressive generative capabilities of these models depend heavily on high-quality image-text datasets, which involve collecting image-text data from the web. This practice raises significant privacy and copyright concerns in the community. The pre-training and fine-tuning processes of text-to-image diffusion models can cause copyright infringement, as they utilize unauthorized datasets published by human artists or stock-image websites [2, 7, 52, 39, 38].
@@
Previous works [4, 11, 23, 13, 12, 32] have attempted membership inference on diffusion models. Carlini et al. [4] employ LiRA (Likelihood Ratio Attack) [3] to perform membership inference on diffusion models. LiRA requires training multiple shadow models to estimate the likelihood ratios of a data point from different models, which incurs high training overhead (e.g., 16 shadow models for DDPM [17] on CIFAR-10 [25]), making it neither scalable nor applicable to text-to-image diffusion models. Other query-based membership inference methods [32, 11, 23, 13] design and compute indicators to evaluate whether a given data point belongs to the member set. These methods require only a few or even a single shadow model, making them scalable to larger text-to-image diffusion models. However, these methods mainly estimate model memorization for data points and do not fully utilize the conditional distribution of image-text pairs. Consequently, they achieve limited success only under excessively high training iterations and fail under real-world training scenarios with common data augmentation methods (Tab. 2), which does not reflect real training scenarios. Text-to-image diffusion models have been demonstrated excellent performance in zero-shot image generation [40, 36, 1], indicating their strong generalization, which makes it difficult to distinguish membership by directly measuring overfitting to data points. And due to the stochasticity of diffusion training loss [17, 40], this kind of measuring becomes more challenging.
@@
To address the challenges, we first identify a Conditional Overfitting phenomenon of text-to-image diffusion models with empirical validation, where the models exhibit more significant overfitting to the conditional distribution of the images given the corresponding text prompts than the marginal distribution of the images only. It inspires the revealing of membership by leveraging the overfitting difference. Based on it, we propose to perform membership inference on text-to-image diffusion models via Conditional Likelihood Discrepancy (CLiD). Specifically, CLiD quantifies overfitting difference analytically by utilizing Kullback-Leibler (KL) divergence as the distance metric and derives a membership inference indicator that estimates the discrepancy between conditional likelihood of image-text pairs and likelihood of images. We approximate the likelihoods by employing Monte Carlo sampling on their ELBOs (Evidence Lower Bounds), and design two membership inference methods: a threshold-based method $\text{CLiD}_{th}$ and a feature vector-based method $\text{CLiD}_{vec}$.
@@
We conduct extensive experiments on three text-to-image datasets [29, 57, 27] with different distributions and sizes using the mainstream open-sourced text-to-image diffusion models [8, 41] under both fine-tuning and pre-training settings. First, our methods consistently outperform existing baselines across various data distributions and training scenarios. For example, $\text{CLiD}_{vec}$ achieves average improvements over the best baseline by 28.7%, 29.8%, and 55.9% in terms of ASR, AUC, and TPR@1%FPR metrics, respectively. Second, we suggest evaluating the membership inference methods by observing the effectiveness trajectory with increasing training steps of a model. The experimental results show our methods reveal membership earlier with fewer training steps. Next, ablation studies further demonstrate the effect of our CLiD indicator, even with fewer query count, our method still outperforms baseline methods (Fig. 3). Last, experiments show that our methods exhibit stronger resistance to data augmentation, and exhibit resistance to even adaptive defenses.
@@ 3.1 Threat Model
We consider a grey-box setting consistent with previous query-based methods [32, 11, 23, 13]. This setting assumes that $\mathcal{A}$ has access to the intermediate outputs of models without knowledge of specific model parameters. For the given image-text data point $(\mathbf{x},\mathbf{c})$, we assume that $\mathbf{x}$ and $\mathbf{c}$ always correspond within the dataset $\mathbb{D}$. This assumption is evident in scenarios where dataset copyright owners perform membership inference to audit usage.
Conversely, challenger $\mathcal{C}$ can mitigate the effectiveness of membership inference during training by utilizing data augmentation or even adaptive defense methods, which we discuss in Sec. 4.5. Our work primarily focuses on the fine-tuning scenarios of text-to-image diffusion models because the weights of pre-trained models are readily available, making this scenario more prone to copyright and privacy risks [50, 35]. And we also conduct experiments on pre-trained text-to-image diffusion models (Sec. 4.6) to demonstrate the effectiveness of our method even in pre-training scenarios.
@@ 3.2 Conditional Overfitting Phenomenon
The rationale behind previous studies primarily hinges on the overfitting of diffusion models to training data (usually image data $\mathbf{x}$[32, 11, 23]. This overfitting tends to result in lower estimation errors for images in the member set (training data) compared to those in the hold-out set during the diffusion process. Various indicators [32, 11, 23] are designed based on this to expose membership information. Specifically, let $q_{\text{mem}}$ and $q_{\text{out}}$ represent the image distributions of the member set and the hold-out set, respectively. $p$ represents the diffusion’s estimated distribution, and $D$ denotes a distance metric (which will be specified later). This rationale can be formulated as:
@@ 3.3 Condition Likelihood Discrepancy
(Proof in the Appendix B) When using $D=D_{KL}$ as distance metric, Assumption 3.1 is equivalent to:
@@ 3.4 Implementation of CLiD-MI
where $\mathbf{c}_{i}^{*}\in\mathbb{C}$. In subsequent parts, we employ their mean or treat them as feature vectors to expose member information. We will introduce how to obtain $\mathbb{C}$ in Sec. 3.5.
@@ 4.1 Setups
Datasets and models. For the fine-tuning setting, we select $416/417$ samples on Pokémon [27], $/2500$ 2500 samples on MS-COCO [29] and $/10,000$ 10 , 000 samples on Flickr [57] as the member/hold-out dataset, respectively. These three datasets involve diverse data distributions and dataset sizes. We use the most widely used text-to-image diffusion model, Stable Diffusion v1-4111 https://huggingface.co/CompVis/stable-diffusion-v1-4 [8], as the target model to fine-tune it on these three datasets. For the pre-training setting, we utilize LAION dataset [45] to conduct experiments on Stable Diffusion v1-5222https://huggingface.co/runwayml/stable-diffusion-v1-5 [41].
Training setups. Previous membership inference on text-to-image diffusion models usually relies on strong overfitting settings. To evaluate the performance more realistically, we consider the two following setups for fine-tuning: (1) Over-training. Following the previous works [11, 23, 13], we fine-tune 15,000 steps on Pokemon datasets, and 150,000 steps on MS-COCO and Flickr (with only $/2500$ 2500 dataset size). (2) Real-world training. Considering that trainers typically do not train for such high steps, we recalibrate the steps based on the training steps/dataset size ratio (approximately 20) of official fine-tuning scripts on Huggingface [19]. Thus, we train 7,500 steps, 50,000 steps and 200,000 steps for the Pokémon, MS-COCO and Flickr datasets, respectively. Additionally, we employ the default data augmentation (Random-Crop and Random-Flip [20]) in training codes [20] to simulate real-world scenarios.
@@
Implementation details. Our evaluation follows the setup of representative membership inference works [3, 4]. It is important to note that some implementations [21, 24] of previous works assume access to a portion of the exact member set and the hold-out set to obtain a threshold for calculating ASR or to train a classification network [21]. This assumption does not align with real-world scenarios. Therefore, we strictly adhere to the fundamental assumption of membership inference [3, 13]: knowing only the overall dataset without any knowledge of the member set and the hold-out set data points. Hence, we firstly train a shadow model to obtain the $\alpha$ for Eq. (16), classifiers for Eq. (18) and the threshold $\tau$ for calculating ASR with auxiliary datasets of the same distribution. Then we perform the test on the target models. Other implementation details are provided in Appendix D.
@@ 4.2 Main Results
Over-training scenario. In Tab. 1, models are trained for excessive steps on all three datasets, resulting in significant overfitting. We observe that under this over-training scenario, both of our methods nearly achieve ideal binary classification effectiveness. For instance, $\text{CLiD}_{th}$ achieves over 99% ASR, AUC and TPR@1%FPR value on the MS-COCO dataset [29]. With this training setup, the metrics for different baselines are very similar. Even the simplest loss-based method [32] (with the query count of 1) also yields satisfactory results compared with other high query count methods. Therefore, we emphasize that such unrealistic over-training scenario fail to adequately reflect the effectiveness differences among various membership inference methods.
Real-world training scenario. In Tab. 2, we adjust the training steps simulating real-word training scenario [19] and utilize default data augmentation [20]. The best value of ASR and AUC of baseline methods decreases to around 65%, and the best value of TPR@1%FPR decreases to around 5%, indicating insufficient effectiveness of previous member inference methods in real-world training scenarios of text-to-image diffusion models. In contrast, our methods maintain ASR above 86% and AUC above 93%, exceeding the best baseline values by about 30%. The TPR@1%FPR of our methods exceeds the best baseline values by 50%~60%. The results demonstrate the effectiveness of our methods across various data distributions and scales in real-world training scenarios.
@@ 4.3 Performance on Various Training Steps
Figure 2: Effectiveness trajectory on various training steps.
In Fig. 2, we fine-tune Stable Diffusion models with MS-COCO dataset for varying training steps under real-world training setting and report the AUC values of different membership inference methods. It can be observed that as the training progresses, $\text{CLiD}_{th}$ exhibits a significantly faster increase in effectiveness trajectory. By $25,000$ steps, $\text{CLiD}_{th}$ effectively exposes membership information, whereas other baselines achieve similar results only at around $150,000$ steps. This demonstrates that our method can effectively reveal membership information when the overfitting degree of the text-to-image diffusion model is weaker.
@@ 4.4 Ablation Study
Figure 3: Performance of $\text{CLiD}_{th}$ under various Monte Carlo sampling numbers (i.e., query count). The legend labels are sorted in ascending order of AUC values and SecMI. $M/N=0$ indicates that the corresponding term is not utilized in Eq. (16).
To conduct ablation study, we vary the Monte Carlo sampling count in Eq. (14) and Eq. (15), perform $\text{CLiD}_{th}$ with MS-COCO dataset under real-world training setting and report the AUC values in Fig. 3. To further compare the effects of Eq. (14) and Eq. (15), we discard each term in Eq. (16) and denote it as $M/N=0$. We also include the result of the best baseline, SecMI [11], as a comparison.
@@
Monte Carlo sampling numbers. In Fig. 3, we observe that when setting $M=N$, the performance improves as the number of Monte Carlo sampling increases. And the performance is improved slightly when $M,N>3$. Hence, we set $M,N=3$ to ensure the balance between a low query count and satisfied performance. Moreover, the results of "M=1, N=1" and "SecMI" also demonstrate that $\text{CLiD}_{th}$ outperforms previous works even with a fewer query count.
4.5 Resistance to Defense
@@
Table 4: Effectiveness of $\text{CLiD}_{th}$ in adaptive defense. We calculate the FID [15] with $10,000$ unseen MS-COCO samples to access the model utility.
@@
Table 5: The performance of different methods on Stable Diffusion v1-5  [41] in pre-training setting. The best results are highlighted in bold.
Method LAION Query
ASR AUC TPR@1%FPR
Loss 51.14 51.77 0.50 1
PIA 51.32 51.36 0.56 2
M. C. 51.92 52.88 1.11 3
SecMI 53.92 54.81 1.68 12
Fluc 56.76 59.65 1.52 20
$\text{CLiD}_{th}$ 61.32 67.64 3.44 15
4.6 Performance on Pre-trained Models
In this part, we evaluate the effectiveness of membership inference methods for the pre-training setting. To ensure the distribution consistency between the member and hold-out set, we respectively select $2500$ samples from the LAION-Aesthetics v2 5+ and LAION-2B MultiTranslated [45] as member/hold-out set following [12], and conduct membership inference on Stable Diffusion v1-5 [40]. As shown in Tab. 5, our method consistently outperforms the baselines across all three metrics.

### v3

@@ Abstract
Text-to-image diffusion models have achieved tremendous success in the field of controllable image generation, while also coming along with issues of privacy leakage and data copyrights. Membership inference arises in these contexts as a potential auditing method for detecting unauthorized data usage. While some efforts have been made on diffusion models, they are not applicable to text-to-image diffusion models due to the high computation overhead and enhanced generalization capabilities. In this paper, we first identify a conditional overfitting phenomenon in text-to-image diffusion models, indicating that these models tend to overfit the conditional distribution of images given the corresponding text rather than the marginal distribution of images only. Based on this observation, we derive an analytical indicator, namely Conditional Likelihood Discrepancy (CLiD), to perform membership inference, which reduces the stochasticity in estimating memorization of individual samples. Experimental results demonstrate that our method significantly outperforms previous methods across various data distributions and dataset scales. Additionally, our method shows superior resistance to overfitting mitigation strategies, such as early stopping and data augmentation.
@@ 1 Introduction
Text-to-image diffusion models have achieved remarkable success in the guided generation of diverse, high-quality images based on text prompts, such as Stable Diffusion [46, 42], DALLE-2 [43], Imagen [49], and DeepFloyd-IF [31]. These models are increasingly adopted by users to create photorealistic images that align with desired semantics. Moreover, they can generate images of specific concepts [32] or styles [61] when fine-tuned on relevant datasets. However, the impressive generative capabilities of these models depend heavily on high-quality image-text datasets, which involve collecting image-text data from the web. This practice raises significant privacy and copyright concerns in the community [5, 18]. The pretraining and fine-tuning processes of text-to-image diffusion models can cause copyright infringement, as they utilize unauthorized datasets published by human artists or stock-image websites [2, 10, 58, 45, 44].
@@
Previous works [5, 15, 28, 17, 16, 38] have attempted membership inference on diffusion models. Carlini et al. [5] employ LiRA (Likelihood Ratio Attack) [4] to perform membership inference on diffusion models. LiRA requires training multiple shadow models to estimate the likelihood ratios of a data point from different models, which incurs high training overhead (e.g., 16 shadow models for DDPM [22] on CIFAR-10 [30]), making it neither scalable nor applicable to text-to-image diffusion models. Other query-based membership inference methods [38, 15, 28, 17] design and compute indicators to evaluate whether a given data point belongs to the member set. These methods require only a few or even a single shadow model, making them scalable to larger text-to-image diffusion models. However, these methods mainly estimate model memorization for data points and do not fully utilize the conditional distribution of image-text pairs. Consequently, they achieve limited success only under excessively high training steps and fail under real steps or common data augmentation methods (Tab. 2), which do not reflect real training scenarios. Text-to-image diffusion models have demonstrated excellent performance in zero-shot image generation [46, 42, 1], indicating their strong generalization, which makes it difficult to distinguish membership by directly measuring overfitting to data points. And due to the stochasticity of diffusion training loss [22, 46], this kind of measuring becomes more challenging.
@@
To address the challenges, we firstly identify a Conditional Overfitting phenomenon of text-to-image diffusion models with empirical validation, where the models exhibit more significant overfitting to the conditional distribution of the images given the corresponding text than the marginal distribution of the images only. It inspires the revealing of membership by leveraging the overfitting difference. Based on it, we propose to perform membership inference on text-to-image diffusion models via Conditional Likelihood Discrepancy (CLiD). Specifically, CLiD quantifies overfitting difference analytically by utilizing Kullback-Leibler (KL) divergence as the distance metric and derives a membership inference indicator that estimates the discrepancy between the conditional likelihood of image-text pairs and the likelihood of images only. We approximate the likelihoods by employing Monte Carlo sampling on their ELBOs (Evidence Lower Bounds), and design two membership inference methods: a threshold-based method $\text{CLiD}_{th}$ and a feature vector-based method $\text{CLiD}_{vec}$.
@@
We conduct extensive experiments on three text-to-image datasets [35, 66, 32] with various data distributions and dataset scales, using the mainstream open-sourced text-to-image diffusion models [11, 47] under both fine-tuning and pretraining settings. First, our methods consistently outperform existing baselines across various data distributions and training scenarios, including fine-tuning settings and the pretraining setting. Second, our experiments on fine-tuning settings with different training steps (Sec. 4.2) reveal that excessively high step/image ratios cause overfitting, leading to hallucination success; and we develop a more realistic pretraining setting following [16, 13], where our experiments reveal the insufficient effect of existing membership inference works [15, 28, 17, 38]. Third, our comparison experiment with varying training steps (Sec. 4.3) indicates that the effectiveness of membership inference grows with higher step/image ratios and should be evaluated under reasonable settings for realistic results. Next, ablation studies further demonstrate the effect of our CLiD indicator, even with fewer query count, our method still outperforms baseline methods (Fig. 3). Last, experiments show that our methods exhibit stronger resistance to data augmentation, and exhibit resistance to even adaptive defenses.
@@ 3.1 Threat Model
We consider a grey-box setting  [^1] Note that in most real-world scenarios, the requirements for $\mathcal{A}$ in gray-box and white-box settings are nearly identical. We use this terminology here for consistency with previous works [15, 28]. consistent with previous query-based methods [38, 15, 28, 17]. This setting assumes that $\mathcal{A}$ has access to the intermediate outputs of models without knowledge of specific model parameters. For the given image-text data point $(\mathbf{x},\mathbf{c})$, we assume that $\mathbf{x}$ and $\mathbf{c}$ always correspond within the dataset $\mathbb{D}$. This assumption is evident in scenarios where dataset copyright owners perform membership inference to audit usage. And we also consider a weaker assumption of conducting membership inference without the groundtruth text in Sec. 4.6.
Conversely, challenger $\mathcal{C}$ can mitigate the effectiveness of membership inference during training by utilizing data augmentation or even adaptive defense methods, which we discuss in Sec. 4.5. Our work primarily focuses on fine-tuning scenarios because the weights of pretrained models are readily available, making this scenario more prone to copyright risks [56, 41]. Numerous projects are implemented by fine-tuning open-source models on specific datasets [24, 60, 3, 64]. We also conduct experiments on pretrained text-to-image diffusion models (Tab. 2) to demonstrate the effectiveness of our method even in pretraining scenarios.
@@ 3.2 Conditional Overfitting Phenomenon
The rationale behind previous studies primarily hinges on the overfitting of diffusion models to training data (usually image data $\mathbf{x}$[38, 15, 28, 7, 8]. This overfitting tends to result in lower estimation errors for images in the member set (training data) compared to those in the hold-out set during the diffusion process. Various indicators [38, 15, 28] are designed based on this to expose membership information. Specifically, let $q_{\text{mem}}$ and $q_{\text{out}}$ represent the image distributions of the member set and the hold-out set, respectively. $p$ represents the diffusion models’ estimated distribution, and $D$ denotes a distance metric (which will be specified later). This rationale can be formulated as:
@@ 3.3 Condition Likelihood Discrepancy
(Proof in Appendix B) When using $D=D_{KL}$ as distance metric, Assumption 3.1 is equivalent to:
@@ 3.4 Implementation of CLiD-MI
where $\mathbf{c}_{i}^{*}\in\mathbb{C}$. In subsequent parts, we employ their mean or treat them as feature vectors to reveal membership information. We will introduce how to obtain $\mathbb{C}$ in Sec. 3.5.
@@ 4.1 Setups
Datasets and models. For the fine-tuning setting, we select $416/417$ samples on Pokémon [32], $/2500$ 2500 samples on MS-COCO [35] and $/10,000$ 10 , 000 samples on Flickr [66] as the member/hold-out dataset, respectively. These three datasets involve diverse data distributions and dataset scales. We use the most widely used text-to-image diffusion model, Stable Diffusion v1-4222 https://huggingface.co/CompVis/stable-diffusion-v1-4 [11], as the target model to fine-tune it on these three datasets. For the pretraining setting, we conduct experiments on Stable Diffusion v1-5333https://huggingface.co/runwayml/stable-diffusion-v1-5 [47] using the processed LAION dataset [51] (detailed in Sec. 4.2) to minimize distribution shift [16, 13].
Fine-tuning setups. For fine-tuning, previous membership inference on text-to-image diffusion models usually relies on strong overfitting settings. To evaluate the performance more realistically, we consider the two following setups: (1) Over-training. Following the previous works [15, 28, 17], we fine-tune 15,000 steps on Pokemon datasets, and 150,000 steps on MS-COCO and Flickr (with only $/2500$ 2500 dataset size). (2) Real-world training. Considering that trainers typically do not train for such high steps, we recalibrate the steps based on the training steps/dataset size ratio (approximately 20) of official fine-tuning scripts on Huggingface [24]. Thus, we train 7,500 steps, 50,000 steps and 200,000 steps for the Pokémon, MS-COCO and Flickr datasets, respectively. Additionally, we employ the default data augmentation (Random-Crop and Random-Flip [25]) in training codes [25] to simulate real-world scenarios.
@@
Implementation details. Our evaluation follows the setup of representative membership inference works [4, 5]. It is important to note that some implementations [26, 29] of previous works assume access to a portion of the exact member set and the hold-out set to obtain a threshold for calculating ASR or to train a classification network [26]. This assumption does not align with real-world scenarios. Therefore, we strictly adhere to the fundamental assumption of membership inference [4, 17]: knowing only the overall dataset without any knowledge of the member/hold-out split. Hence, we first train a shadow model to obtain the $\alpha$ for Eq. (16), classifiers for Eq. (18) and the threshold $\tau$ for calculating ASR with auxiliary datasets of the same distribution. Then we perform the test on the target models. Other implementation details are provided in Appendix D.
@@ 4.2 Main Results
Over-training setting (fine-tuning). In Tab. 1, models are trained for excessive steps on all three datasets, resulting in significant overfitting. We observe that under this over-training scenario, both of our methods nearly achieve ideal binary classification effectiveness. For instance, $\text{CLiD}_{th}$ achieves over 99% ASR, AUC and TPR@1%FPR value on the MS-COCO dataset [35]. With this training setup, the metrics for different baselines are very similar. Even the simplest loss-based method [38] (with the query count of 1) also yields satisfactory results compared with other high query count methods. Therefore, we emphasize: This unrealistic over-training setting fails to adequately reflect the effectiveness differences among various membership inference methods.
Real-world training setting (fine-tuning). In Tab. 2, we adjust the training steps simulating real-word training scenario [24] and utilize default data augmentation [25]. The best value of ASR and AUC of baseline methods decreases to around 65%, and the best value of TPR@1%FPR decreases to around 5%, indicating insufficient effectiveness of previous member inference methods in real-world training scenarios of text-to-image diffusion models. In contrast, our methods maintain ASR above 86% and AUC above 93%, exceeding the best baseline values by about 30%. The TPR@1%FPR of our methods exceeds the best baseline values by 50%~60%. The results demonstrate the effectiveness of our methods across various data distributions and scales in real-world training scenarios.
Pretraining setting. For the pretraining setting, we adopt a stringent and realistic membership inference setting based on previous works [16, 13]. (1) To ensure the distribution consistency between the member and hold-out set, we respectively select $2500$ samples from the LAION-Aesthetics v2 5+ and LAION-2B MultiTranslated [51] as member/hold-out set following [16]; (2) We filter out samples containing non-English characters to ensure there are no other "distinguishable tails" [13] in the dataset [^4] Das et al. [13] indicates that MultiTranslated-LAION dataset contains fewer non-English characters than the LAION dataset due to the use of the translation model. . We conduct membership inference on Stable Diffusion v1-5 [46]. As shown in Tab. 2, our method consistently outperforms the baselines across all three metrics.
Method LAION Query
ASR AUC TPR@1%FPR
Loss 51.78 50.90 1.75 1
PIA 52.13 52.42 1.25 2
M. C. 53.18 53.96 1.25 3
SecMI 57.43 58.59 2.45 12
PFAMI 59.08 61.11 1.45 20
$\text{CLiD}_{th}$ 64.53 67.82 5.01 15
Table 3: The performance of membership inference methods on Stable Diffusion v1-5 [47] in pretraining setting. We utilize the processed LAION dataset to ensure the distribution consistency between member / hold-out sets [16, 13]. The best results are highlighted in bold.
Figure 2: Effectiveness trajectory on various training steps.
@@ 4.3 Performance on Various Training Steps
To explore this, we fine-tune Stable Diffusion models with the MS-COCO dataset for varying training steps under real-world training setting and report the AUC values of different membership inference methods in Fig. 2. It can be observed that as the training progresses, $\text{CLiD}_{th}$ exhibits a significantly faster increase in effectiveness trajectory. By $25,000$ steps, $\text{CLiD}_{th}$ effectively exposes membership information, whereas other baselines achieve similar results only at around $150,000$ steps. This demonstrates that our method can effectively reveal membership information when the overfitting degree of the text-to-image diffusion model is much weaker.
@@ 4.4 Ablation Study
Figure 3: Performance of $\text{CLiD}_{th}$ and SecMI under various Monte Carlo sampling numbers (i.e., query count). The legend labels are sorted in ascending order by AUC values.
To conduct an ablation study, we vary the Monte Carlo sampling count in Eq. (14) and Eq. (15), perform $\text{CLiD}_{th}$ with MS-COCO dataset under real-world training setting and report the AUC values in Fig. 3. To further compare the effects of Eq. (14) and Eq. (15), we discard each term in Eq. (16) and denote it as $M/N=0$. We also include the result of the best baseline, SecMI [15], as a comparison.
@@
Monte Carlo sampling numbers. In Fig. 3, we observe that when setting $M=N$, the performance improves as the number of Monte Carlo sampling increases. And the performance is improved slightly when $M,N>3$. Hence, we set $M,N=3$ to ensure the balance between a low query count and satisfied performance. Moreover, the experiment results of "M=1, N=1" and "SecMI" also demonstrate: $\text{CLiD}_{th}$ outperforms previous works even with a much fewer query count.
@@
Table 5: Effectiveness of $\text{CLiD}_{th}$ in adaptive defense. We calculate the FID [20] with $10,000$ unseen MS-COCO samples to assess the model utility.
@@
4.5 Resistance to Defense
4.6 Weaker Assumption
Although in Sec. 3.1 we assume that the adversary can access the entire image-text pairs based on the real-world data usage auditing scenario, we also consider a weaker assumption: the adversary can only access the image without the corresponding text.
In this scenario, we first generate pseudo-text corresponding to the images using an image captioning model (BLIP [34] in our experiments), and then conduct CLiD-MI based on the image-pseudo_text pairs. In Tab. 6, we observe that our method still broadly outperforms baselines. We believe this is because the pseudo-text preserves the image’s key semantics, keeping our methods effective.
Table 6: Results without access to the corresponding text under Over-training setting and Real-world training setting. We fine-tune MS-COCO on SDv1-4. Key results are highlighted as Tab. 1.
Method Over-training (Pseudo-Text) Real-world training (Pseudo-Text) Query
ASR AUC TPR@1%FPR ASR AUC TPR@1%FPR
Loss 73.80 81.01 9.71 56.08 58.47 1.60 1
PIA 61.40 65.75 1.20 53.44 54.38 1.52 2
M. C. 74.36 81.55 11.28 56.68 60.00 1.28 3
SecMI 82.04 88.97 40.80 60.48 64.04 3.28 12
PFAMI 91.56 95.16 68.16 58.12 59.77 2.64 20
$\text{CLiD}_{th}$ 92.84 95.43 72.36 76.16 83.27 19.76 15
$\text{CLiD}_{vec}$ 93.26 96.59 71.73 77.76 84.48 18.06 15